A gun of mass 2.0kg fired a bullet of mass 0.02kg wit d velocity of 15o/m, calculate d kinetic energy of d gun and kinetic energy of d bullet. Formula =1/2x mass x velocity.
1
Expert's answer
2020-01-28T13:13:59-0500

Given:

"m_1=2.0\:\rm kg;"

"m_2=0.02\:\rm kg;"

"v_2=150\:\rm m/s;"

The speed of the gun after shot


"v_1=\frac{m_2v_2}{m_1}=\frac{0.02\times 150}{2.0}=1.5\:\rm m/s"

Hence, the kinetic energy of the gun:


"E_1=\frac{m_1v_1^2}{2}=\frac{2.0\times 1.5^2}{2}=2.25\:\rm J"

The kinetic energy of the bullet:

"E_2=\frac{m_2v_2^2}{2}=\frac{0.02\times 150^2}{2}=225\:\rm J"